OpenStudy (anonymous):

Solve x divided by nine is less than or equal to negative three for x. Show your work.

OpenStudy (***[isuru]***):

\[\frac{ x }{ 9 } \le -3\] multiply both sides by 9 \[\frac{ x }{ 9 }\times 9 \le -3 \times 9\] which will give u \[x \le -27\] got it :) ?

OpenStudy (anonymous):

THX SOOO MUCH!!! :DDD!!!

OpenStudy (***[isuru]***):

u r welcome!